(x^3-3x^2-108x)/(x-12)
Answer by venugopalramana(3286)

(x^3-3x^2-108x)/(x-12)
=X(X^2-3X-108)/(X-12)
=X(X^2-12X+9X-108)/(X-12)
=X{X(X-12)+9(X-12)}/(X-12)
=X(X+9)(X-12)/(X-12)
=X(X+9)